# Augmented Object Intelligence: Making the Analog World Interactable with XR-Objects

### v4

@@ 4.6.1. Time
On average, participants using XR-Objects  required significantly less time (M=217.5 s, SD=58 s) to complete all tasks when compared to the participants using Chatbot (M=286.3, SD=71 seconds), illustrated in Figure 7 and confirmed by a paired t-test (t=-2.8, df=5, p=0.01). This translates to a roughly 31% in task completion time on average for Chatbot users compared to XR-Objects.
[…]
@@ 4.6.3. Form Factor (Phone or HMD)
Figure 9. Marimekko-chart mosaic with the form factor survey results.
[…]
The mosaic chart shows a preference for XR-Objects over Chatbot on the HMD condition.
@@ 5. Applications
Through AOI, we envision XR-Objects  to be useful across a variety of real-world applications. By enabling in situ digital interactions with non-instrumented analog objects, we can expand their utility (e.g., enabling a pot to double as a cooking timer), better synthesize their relevant information (e.g., comparing nutritional value), and overall enable richer interactivity and flexibility in everyday interactions. Here we present five example application scenarios from a broad application space we envision (see Figure 10) that highlight the value of XR-Objects.
Figure 10. Envisioned application space for XR-Objects.
@@ 5.1. Discovery
XR-Objects  enables users to discover new information about their surroundings by simply pointing their device at objects of interest. As shown in Figure 11, a user points their device at a vase containing different flowers and instantly receive information about each flower type, including its name, price, and care instructions. This on-demand, spatial discovery transforms everyday objects into avenues for understanding or appreciation of everyday objects that normally go unnoticed.
Figure 11. Example of discovery through MLLM interactions on top of XR-Objects.
@@ 5.2. Productivity
For productivity, XR-Objects  could enhance physical documents with digital capabilities such as information retrieval and content anchoring. In Figure 12, a user reading a text book asks how it can be used to solve a particular type of equation, and anchors the response to the textbook for future reference. With added capabilities such as real-time optical character recognition (OCR) to digitize text, users could store and share digital copies of their physical documents for versioning and collaboration.
Figure 12. Interaction mock up of using XR-Objects  for productivity on an HMD with direct touch interactions enabled.
@@ 5.3. Learning
XR-Objects  offers immersive and interactive learning experiences by augmenting physical objects with contextual educational content. By pointing their device at an object, users can access relevant explanations, demonstrations, and quizzes that enhance their understanding of the subject matter.
As illustrated in Figure 13, XR-Objects  can facilitate learning about healthy eating habits for children. A child can point their device at a fruit bowl and instantly see information about the different fruits, such as their names, nutritional values, and the vitamins they contain. This interactive learning experience can help children develop a deeper understanding of the importance of a balanced diet and the benefits of various fruits.
Figure 13. Example of anchoring contextual educational content over real-world objects using XR-Objects.
@@ 5.4. IoT Connectivity
XR-Objects’s XR interface is complemented by a MLLM backend, unlike traditional IoT control interfaces that often limit interactions with devices to discrete apps. Through its object tracking, XR-Objects  enables users to interact with their IoT devices in a spatial context, allowing for real-time visual feedback and control within their immediate environment.
Figure 14. XR-Objects  detects devices and provides custom UI elements for IoT control.
@@
5.5. Cooking
Unlike traditional cooking aids, which rely on static recipes or digital screens detached from the cooking environment, XR-Objects  integrates digital intelligence directly into the kitchen, making the cooking process informative and engaging (Figure 1). In our augmented cooking app, as the user places ingredients on the kitchen counter, our system recognizes each item and projects relevant information, such as freshness indicators, nutritional facts, and potential allergens, directly on the ingredients. Users can interact through voice commands or touch elements to ask about potential recipes or to compare ingredients. Using stopwatch or countdown timers, the system embeds the guidance into the cooking space itself. The user can share the final product with their contacts.
@@ 6. Discussion
Our daily environment is ubiquitously augmented with various forms of annotations, from product packaging and price tags to traffic signs and personalized post-it reminders. This ubiquitous augmentation, a rudimentary form of augmented reality, is facilitated by the advent of language, writing, and scalable printing technologies. It serves as a means to asynchronously convey contextual information, albeit in a static and limited manner, requiring manual interpretation and application by the user. Although machine-readable markers like barcodes, QR codes, and NFC tags have simplified certain interactions, they fall short in offering dynamic, object-relevant actionable insights.
@@
Looking ahead, XR-Objects  could evolve to become even more attuned to the user’s context (e.g. current location and activity, object relationships, persistence of object and data, see Figure 10) to further customize actions and information based on historical interactions with objects in specific settings. For instance, viewing a food item’s packaging in a store might trigger suggestions for understanding its allergens and nutritional content, while the same item at home could offer cooking instructions and allow the user to directly set a digital cooking timer to the appropriate duration. On certain smartphones, for instance, Siri Suggestions [^10] Siri Suggestions on iPhone: https://support.apple.com/guide/iphone/about-siri-suggestions-iph6f94af287/ios already offer context-aware recommendations. The functionality of this sort of context-awareness could be extended using AR interfaces.
@@
The integration of emerging AGIs (Morris et al., 2024), exemplified by models like Gemini or GPT-4, opens up new opportunities for autonomous problem-solving within XR environments. AGI’s potential to dynamically generate user interfaces and visualizations in response to user queries could transform the way we interact with our physical world. For instance, the user could ask the system to visualize the nutritional values of a product in a pie chart, i.e., by generating the code to create the user interface and graphs on-the-fly without being pre-programmed to create the chart. Going a step further, one may imagine AGI-driven systems that not only respond to user prompts but also proactively offer assistance surfaced through a new action in the context menu, such as assembling a set of Lego blocks into a desired structure through real-time, augmented instructions.
@@
As we adopt these new innovations in AI and XR, we are likely to see significant changes in how we interact with the physical objects around us. We envision a future where physical items no longer need conventional labels or tags, relying instead on AOI for context and interaction. The merging of digital and physical worlds might bring about new connections, like direct links between digital files and physical items. This could start a new phase where the digital and physical worlds blend together smoothly, without clear boundaries.
@@ 7. Conclusion
In this paper, we introduced Augmented Object Intelligence  (AOI), a novel paradigm that seamlessly integrates digital capabilities into physical objects through the use of XR-Objects. Our prototype system demonstrates the potential of AOI to transform how users interact with their surroundings by leveraging advancements in computer vision, spatial understanding, and Multimodal -LLM. The results of our user study show that XR-Objects  significantly outperforms traditional multimodal AI interfaces, with participants completing tasks an average of 24% shorter time and reporting higher levels of satisfaction, ease of use, and perceived responsiveness. By enabling familiar, interactions with everyday objects through anchored AR content and natural language processing, XR-Objects  paves the way for a future where the boundaries between the physical and digital worlds become increasingly blurred. As we continue to expand the capabilities of XR-Objects, we envision a wide range of applications spanning domains such as cooking, productivity, and IoT connectivity, ultimately leading to a more engaging, efficient, and immersive way of interacting with our surroundings.
@@ A.2. Form Factor Survey
For each question, participants selected either Chatbot or XR-Objects. They did this questionnaire twice, once we were asking participants to think XR-Objectswas going to run on a Phone, the second time thinking XR-Objectswould run on a headset form factor. This questionnaire included the adapted HALIE questions (H, J, S, R, ER, E2, EN) as well as an additional set of questions detailed below:

### v5

@@ 4.2.1. Time
On average, participants using XR-Objects  required significantly less time (M=217.5s, SD=58s) to complete all tasks when compared to the participants using Chatbot (M=286.3s, SD=71s), illustrated in Figure 14 and confirmed by a paired t-test (t=-2.8, df=5, p=0.01). This translates to a roughly 31% in task completion time on average for Chatbot users compared to XR-Objects.
@@ 4.2.3. Form Factor (Phone or HMD)
4.3. Qualitative Feedback
We provide key insights from the responses that participants provided in their completed surveys.
Helpfulness & Efficiency. Users valued the system’s streamlined interactions: ”It saves me from looking up info myself… I just ask and it finds the info for me” (P1). ”Object selection and comparison was very intuitive; it was easy to get information in exactly the context I needed” (P3).
Comparative Advantage. The direct comparison with existing solutions like Google Lens and traditional LLMs was enlightening: ”I was able to complete the same tasks much faster + easier” (P1). ”This has a lot more options, and is more flexible [in] what information it can provide” (P2). ”Comparing products is very helpful… less wordy than [Chatbot] and gives an answer” (P2).
Possibilities for UX Improvement. Several participants pointed out ergonomic challenges, e.g., the need to hold the phone at eye level, which informs future glass interactions as discussed in Section 6: ”have to raise the phone at eye level, which is tiring” (P1).
@@ 5. Applications
Through AOI, we envision XR-Objects  to be useful across a variety of real-world applications. By enabling in situ digital interactions with non-instrumented analog objects, we can expand their utility (e.g., enabling a pot to double as a cooking timer), better synthesize their relevant information (e.g., comparing nutritional value), and overall enable richer interactivity and flexibility in everyday interactions. Next, we present five example application scenarios from a broad design space we envision as illustrated in Figure 8, highlighting the value of XR-Objects  and its potential use cases.
Figure 8. Envisioned design space for XR-Objects  use cases.
@@
5.1. Cooking
Unlike traditional cooking aids, which rely on static recipes or digital screens detached from the cooking environment, XR-Objects  integrates digital intelligence directly into the kitchen, making the cooking process informative and engaging (Figure 1). In our augmented cooking app, as the user places ingredients on the kitchen counter, our system recognizes each item and projects relevant information, such as nutritional facts, potential allergens, or freshness indicators, directly on the ingredients. Users can interact through voice commands or touch elements to ask about potential recipes or to compare ingredients. Using stopwatch or countdown timers, the system embeds the guidance into the cooking space itself. The user can further share the final product with their contacts.
This scenario shows how XR-Objects  can assist users with multi-step tasks. For instance, P5 noted: “XR-Objects’s main benefit compared to Lens is spatializing output, [which] is super helpful for labeling multi-step tasks like cooking or mechanical fixing.” While we illustrated cooking in this example, we envision our system can scale to further multi-step use cases, such as mechanical fixing.
Figure 9. Real-time spatial assistance in selecting the appropriate laundry detergent at a store.
[…]
Left: detergents on a shelf. Right
5.2. Shopping
As discussed in Section 4, XR-Objects  can serve as an assistant when browsing and comparing items. For instance, at a store, a user might want to get further information about a product, such as the unit price, calorie information, or cheaper alternatives. Figure 9 shows an example where a tourist asks which of the detergents is suitable for black clothes, as the product information is written in a foreign language. In the future, these experiences could be further personalized by, e.g., automatically filtering all the products identified in the scene based on the user’s personal profile and recommending the one that best suits their needs (Xu et al., 2022).
@@ 5.3. Discovery
XR-Objects  enables users to discover new information about their surroundings by simply pointing their device at objects of interest. As shown in Figure 10, a user points their device at a vase containing different flowers and instantly receives information about each flower type, including its name, average price, or care instructions. This on-demand, spatial discovery could transform everyday objects that typically go unnoticed into avenues for appreciation.
Figure 10. Discovering plant species in the environment through spatial, on-demand MLLM queries.
@@ 5.4. Productivity
Figure 11. Interaction mock-up of using XR-Objects  for productivity on an HMD with direct touch interactions enabled.
@@
For productivity, we envision XR-Objects  could enhance physical documents with digital capabilities such as information retrieval and content anchoring. In Figure 11, a user reading a text book asks how it can be used to solve a particular type of equation, and anchors the response to the textbook for future reference. With further capabilities such as real-time optical character recognition (OCR) to digitize text, users could store and share digital copies of their physical documents for versioning and collaboration.
@@ 5.5. Learning
XR-Objects  can offer immersive or interactive learning experiences by augmenting physical objects with contextual educational content. By pointing their device at an object, users may access relevant explanations or demonstrations that enhance their understanding of the subject matter. For instance, as illustrated in Figure 12, XR-Objects  could facilitate learning healthy eating habits for children. A child can point their device at a fruit bowl and instantly see information about the different fruits, such as their names, nutritional values, and the vitamins they contain. We also envision users leaving spatial notes on objects, and these could have changeable different visibility options. For instance, a user might leave a personal note about something they found out which they would like to attach as a reminder. They might further set the visibility to a specific group, e.g., their family members or coworkers, so that others can see and be aware of the new information.
Figure 12. Example of anchoring contextual educational notes over real-world objects using XR-Objects.
@@ 5.6. IoT Connectivity
XR-Objects’s XR interface is complemented by a MLLM backend, unlike traditional IoT control interfaces that often limit interactions with devices to discrete apps. Through its object tracking, XR-Objects  could enable users to interact with their IoT devices in a spatial context to allow for real-time visual feedback and control within their immediate environment.
Figure 13. XR-Objects  detects devices, such as speakers, and provides custom UI elements for IoT control.
@@ 6. Discussion
Our daily environment is ubiquitously augmented with various forms of annotations, from product packaging and price tags to traffic signs and personali post-it reminders. This ubiquitous augmentation, a rudimentary form of AR, is facilitated by language, writing, and scalable printing technologies. It serves as a means to asynchronously convey contextual information, albeit in a static and limited manner, requiring manual interpretation and application by the user. Although machine-readable markers like barcodes, QR codes, and NFC tags have simplified certain interactions, they fall short in offering dynamic, object-relevant actionable insights.
@@
While the list of actions we currently provide in our system is not exhaustive, we view it as a foundation that can be expanded by the community. Looking ahead, we anticipate contributions from researchers and practitioners, e.g., integrating methods such as OmniActions (Li et al., 2024) for custom action suggestions, WorldGaze (Mayer et al., 2020) for gaze input, and GazePointAR (Lee et al., 2024) for pronoun disambiguation in speech interactions. Using these extensions, XR-Objects  could evolve to become even more attuned to the user’s context (e.g. current location and activity, object relationships, persistence of object and data, see Figure 8) to further customize actions and information based on historical interactions with objects in specific settings. For instance, viewing a food item’s packaging in a store might trigger suggestions for understanding its allergens and nutritional content, while the same item at home could offer cooking instructions and allow the user to directly set a digital cooking timer to the appropriate duration. On certain smartphones, for instance, Siri Suggestions [^7] Siri Suggestions on iPhone: https://support.apple.com/guide/iphone/about-siri-suggestions-iph6f94af287/ios already offer context-aware recommendations. The functionality of this sort of context-awareness could be extended using AR interfaces.
XR-Objects  on AR Headsets and Glasses
Participants expressed enthusiasm for XR-Objects  on wearable AR. P3 mentioned the current challenges of typing on headsets and the advantages of voice and selection-based interfaces such as XR-Objects. P2 and P8 emphasized the practicality of using features like spatial timers on glasses for more glanceable interactions. They noted XR-Objects  enhance the ease of comparing and selecting objects. P8 suggested potential cross-device interactions, such as starting a task on glasses and continuing on a phone for detailed comparisons.
As noted in Section 3, a possible implementation on today’s headsets is currently challenging due to restricted access to camera streams (Quest 3) or limited FOV (HoloLens). As an alternative, future projects might explore attaching a webcam onto a Quest 3 and manually calibrating the camera for XR-Objects  detection and interactions. Our open-source repository (Google, 2024b) contains a Unity guide for processing a webcam feed on headsets to implement this. We envision that on headsets, developers could use mid-air gestures, rays, or other hand-based interaction methods (Bawa, 2022).
LLM: Hallucinations and AI Improvements
We recognize LLM as an emerging technology. In this work, our primary focus is enhancing spatial interaction, not solely LLM accuracy. As AI evolves, so will our system, incorporating user feedback to refine LLM interactions based on object-centric data.
In our study, participants appreciated XR-Objects’ reduced risk of hallucination due to its object-instance-based spatialization. P4 noted: “The control and feedback for intermediate steps – which object was recognized, what the model thinks it is – afforded by XR-Objects  provides more confidence that the model isn’t hallucinating and makes it easier to spot when it is.
@@
The integration of emerging AGIs (Morris et al., 2024), foreshadowed by models like Gemini or GPT-4, opens up new opportunities for autonomous problem-solving within XR environments. AGI’s potential to dynamically generate user interfaces in response to user queries could transform the way we interact with our physical world. For instance, the user could ask the system to visualize the nutritional values of a product in a pie chart, i.e., by generating the code to create the user interface and graphs on-the-fly without being pre-programmed to create the chart. Going a step further, one may imagine AGI-driven systems that not only respond to user prompts but also proactively offer assistance surfaced through a new action in the context menu, such as assembling a set of Lego blocks into a desired structure through real-time, augmented instructions.
Object Detection and Selection
Our real-time object detection operates at 31 fps on a Galaxy S21 smartphone without needing pre-processing. MLLM queries respond in approximately three seconds. Currently, the object selection is bound to the MediaPipe classifier’s output. We plan to explore enhancements like subregion selection, as P1 suggested: ”would be nice [to] manipulate the scene with touch gestures [to] lead the AI to detect a thin object.
@@
As we adopt these new innovations in AI and XR, we are likely to see significant changes in how we interact with the physical objects around us (Hirzle et al., 2023; Nebeling et al., 2024). We envision a future where physical items no longer need conventional labels or tags, relying instead on AOI for context and interaction. The merging of digital and physical worlds might bring about new connections, like direct links between digital files and physical items. This could start a new paradigm where the digital and physical worlds blend together smoothly, without clear boundaries.
@@ 7. Conclusion
In this paper, we introduced Augmented Object Intelligence  (AOI), a novel paradigm that seamlessly integrates digital capabilities into physical objects through the use of XR-Objects. Our prototype system demonstrates the potential of AOI to transform how users interact with their surroundings by leveraging advancements in computer vision, spatial understanding, and Multimodal -LLM. The results of our user study show that XR-Objects  significantly outperforms traditional multimodal AI interfaces, with participants completing tasks an average of 24% shorter time and reporting higher levels of satisfaction, ease of use, and perceived responsiveness. By enabling familiar, interactions with everyday objects through anchored AR content and natural language processing, XR-Objects  paves the way for a future where the boundaries between the physical and digital worlds become increasingly blurred. As we continue to expand the capabilities of XR-Objects, we envision a wide range of applications spanning domains such as cooking, productivity, and connectivity, ultimately leading to a more engaging, efficient, and immersive way of interacting with our surroundings.
@@ A.2. Form Factor Survey
For each question, participants selected either Chatbot or XR-Objects. They did this questionnaire twice, once we were asking participants to think XR-Objects  were going to run on a Phone, the second time thinking XR-Objects  would run on a headset form factor. This questionnaire included the adapted HALIE questions (H, J, S, R, ER, E2, EN) as well as an additional set of questions detailed below:
@@
A.3. Results
The analysis of completion time and form factory survey is depicted in Figure 14 and Figure 15, respectively.
[…]
Figure 15. Marimekko-chart mosaic with the form factor survey results, showing a preference for XR-Objects  over Chatbot on the HMD condition.
[…]
The mosaic chart shows a preference for XR-Objects over Chatbot on the HMD condition, compared to the phone condition, where the participants were indifferent across most of the conditions.